Write the slope-intercept form of the equation. (-1,-8) and (-1,9)
Found 2 solutions by rfer, checkley77:
Answer by rfer(16322)

this is an undefined slope
no y int

Answer by checkley77(12844)

(-1,-8) and (-1,9)
Y=mX+b is the line equation when m=slope & b=the y intercept.
m=(y2-y1)/(x2-x1)
m=(9+8)/(-1+1)=17/0 this indicates that the slope is a verticle line through
x=-1 is the line equation.
There is no y intercept.